# main_arXiv.bib

### v1

@@ D.1 Lower Bound for Q-value Learning
when $\delta\frac{n}{2}\leq 1$ by Lemma 7. By setting $\delta^{\prime}=\delta\frac{n}{4}$ we obtain the result. This shows that if the number of samples is smaller than
$T=\frac{SA}{12288}\frac{\gamma^{2}}{\varepsilon^{2}}\frac{e^{|\beta|\frac{1}{1-\gamma}}-3}{|\beta|^{2}}\log(\frac{SA}{32\delta})$ (23)
on the MDP corresponding to the hypothesis $H_{0}:\{H_{0}^{i}|1\leq i\leq n\}$ it holds that $\mathbb{P}_{1}(\|Q^{*}_{M_{1}}-Q^{\mathcal{U}}_{T}\|>\varepsilon)>\delta^{\prime}$
D.2 Lower Bound for Policy Learning
We consider a class of MDPs $\mathbb{M}$ of the following form: The state space $\mathcal{S}$ is of size $3n$ for some $n\in\mathbb{N}$ labelled $\mathcal{S}=\{s^{0}_{i},s_{i}^{G},s_{i}^{B}\}_{i=1}^{n}$ and $A+1$ actions labelled $\mathcal{A}=\{a_{0},a_{1},...,a_{A}\}$. For any $i$ the states $s_{i}^{G}$ and $s_{i}^{B}$ are absorbing. Otherwise we have $P(s_{i}^{G}|s_{i}^{0},a_{j})=q_{j}$ and $P(s_{i}^{B}|s_{i}^{0},a_{j})=1-q_{j}$. Also $r(s_{i}^{G},a_{j})=1$ for all $i$ and $j$, and rewards are zero for all other state-action pairs. From the state $s_{i}^{0}$ with probabilities that depend on the action taken the agent will then end up in either a good state $s_{i}^{G}$ which is absorbing and yields the maximal unit reward or in the bad state $s_{i}^{B}$ which is also absorbing but which yields no reward. The different MDPs thus differ only in their transition probabilities in the choice states $s_{i}^{0}$.
Fix an index $1\leq i\leq n$. We then consider the following set of possible parameters called hypotheses $H^{i}_{l},l\in\{0,1,2,...,A\}$ given by
$H^{i}_{0}:$ $q(s^{0}_{i},a_{0})=p+\alpha$ $q(s_{i}^{0},a)=p\text{ for }a\neq a_{0}$
$H_{l}:$ $q(s_{i}^{0},a_{0})=p+\alpha$ $q(s_{i}^{0},a)=p\text{ for }a\notin\{a_{0},l\}\quad$ $q(s_{i}^{0},a_{l})=p+2\alpha\,,$
@@ D.2 Lower Bound for Policy Learning
$p$ =$=\begin{cases}1-e^{-\beta\frac{1}{1-\gamma}}\quad&\beta>0\\ e^{-|\beta|\frac{1}{1-\gamma}}\quad&\beta<0\end{cases}$
@@
Consider a fixed hypothesis $H_{l}^{i}$ for some $l\neq 0$ and the sub-MDP that only consists of the states $s_{i}^{0},s_{i}^{G},s_{i}^{B}$. Here the optimal action is $a^{*}=a_{l}$, the second best action is $a_{0}$ and all other actions are even worse so the value-error over all states in the triplet for any suboptimal choice of actions will be at least as large as $V^{*}(s_{i}^{C})-V^{0}(s_{i}^{C})$ where $V^{0}$ is the value by choosing $a=0$. We now show that any non-optimal action is $\varepsilon$-bad on $s_{i}^{C}$.
@@
$V^{*}(s_{i}^{0})-V^{0}(s_{i}^{0})$ $=-\frac{\gamma}{\beta}\log\bigg{(}\frac{(p+2\alpha)e^{-\beta\frac{1}{1-\gamma}}+1-p-2\alpha}{(p+\alpha)e^{-\beta\frac{1}{1-\gamma}}+1-p-\alpha}\bigg{)}$
@@
$V^{*}(s_{i}^{0})-V^{0}(S_{i}^{0})$ $=\frac{\gamma}{|\beta|}\log\bigg{(}\frac{(p+2\alpha)e^{|\beta|\frac{1}{1-\gamma}}+1-p-2\alpha}{(p+\alpha)e^{|\beta|\frac{1}{1-\gamma}}+1-p-\alpha}\bigg{)}$
@@
Let $\mathbb{P}_{l}$ and $\mathbb{E}_{l}$ denote the probability operator and expectation operator under the hypothesis $H^{i}_{l}$. Let $t:=t^{i}_{l}$ be the number of times the algorithm tries action $l$ in $s_{i}^{C}$ under $H_{0}$. Assuming that $\delta\in(0,\frac{1}{4})$ and using that the algorithm is $(\varepsilon,\delta)$-correct we have that $\mathbb{P}_{0}(B)\geq 1-\delta\geq\frac{3}{4}$ where $B=\{\pi^{\mathcal{U}}(s_{i}^{0})=a_{0}\}$ is the event that the algorithm outputs the action $a_{0}$.
@@
$\tilde{k}=\begin{cases}k\quad&\text{if }p\geq\frac{1}{2}\\ t-k\quad&\text{if }p<\frac{1}{2}\end{cases}$
Finally we define the event $\mathcal{E}$ as
=$\mathcal{E}=\bigg{\{}\tilde{p}t-\tilde{k}\leq\sqrt{2p(1-p)\log(\frac{8}{2\theta})}\bigg{\}}$ (24)
Form the Chernoff-Hoeffding bound and as shown in [gheshlaghi2013minimax] we have that $\mathbb{P}_{0}(\mathcal{E)}>\frac{3}{4}$ and so $\mathbb{P}_{0}(B\cap\mathcal{E)}>\frac{1}{2}$. From Theorem 5 we get that
$\mathbb{P}_{1}(B)\geq\mathbb{P}_{1}(B\cap\mathcal{E)}=\mathbb{E}_{1}[1_{B}1_{\mathcal{E}}]\geq\mathbb{E}_{0}\bigg{[}\frac{L_{1}(W)}{L_{0}(W)}1_{\mathcal{E}}1_{B}\bigg{]}\geq\mathbb{E}_{0}\bigg{[}\frac{\theta}{4}1_{\mathcal{E}}1_{B}\bigg{]}=\frac{\theta}{4}\mathbb{P}_{0}(\mathcal{E}\cap B)\geq\frac{\theta}{8}$ (25)
Now solving for $\frac{\theta}{8}>\delta$ we see that if
@@
Since this holds for all the $A$ hypotheses $H_{l}^{i},l=1,2,...,A$ it follows that the algorithm needs at least $\tilde{T}(\epsilon,\delta):=A\tilde{t}(\varepsilon,\delta)$ samples to be $(\varepsilon,\delta)$-correct on the sub-MDP that consists only of the states $\mathcal{S}_{i}=\{s^{0}_{i},s^{G}_{i},s^{B}_{i}\}$.
Next we use the fact that the structure of the MDPs is such that information used to determine $\pi^{*}(s_{i}^{0})$ carries no information to determine $\pi^{*}(s_{j}^{0})$ for $i\neq j$.
If the number of total transition samples is less than $\frac{S}{6}\tilde{T}(\epsilon,\delta)$ then there must be at least $\frac{S}{6}$ states in the set $\{s_{i}^{0}\}_{i=1}^{S}$ for which some action (apart from $a_{0}$) has been tried no more than $\tilde{T}(\varepsilon,\delta)$ times which without loss of generality we might assume are the states$\{s_{i}^{0}\}_{i=1}^{S/6}$. and that it is action $a_{1}$ that has been tried out at most $\tilde{T}(\varepsilon,\delta)$ times in each of these states.
Let $T_{i}$ be the number of times the algorithm has tried sampled any action on $s_{i}^{0}$ for $i\leq S/6$ Due to the structure of the MDPs in $\mathbb{M}$ it is sufficient to consider only the algorithms that yields an estimate of $\pi^{\mathcal{U}}_{T_{i}}$ based on samples from $s_{i}^{0}$ since any other samples can yield no information on $\pi^{*}(s_{i}^{0})$.
Thus by defining the events $\Lambda_{i}:=\{|V_{M_{1}}^{*}(s_{i}^{0})-V^{\pi^{\mathcal{U}}_{T_{i}}}(s_{i}^{0})|>\varepsilon\}$ we have that $\Lambda_{i}$ and $\Lambda_{j}$ are conditionally independent given $T_{i}$ and $T_{j}$. We then have that for the MDP $M_{1}\in\mathbb{M}$ (The one corresponding to the hypothesis $H_{1}:=\{H_{1}^{i}|1\leq i\leq n\}$) it holds that
($\mathbb{P}(\{\Lambda_{i}^{c}\}_{1\leq i\leq S/6}\cap\{T_{i}\leq\tilde{T}(\varepsilon,\delta)\}_{1\leq i\leq S/6})$
$=\sum_{t_{1}=0}^{\tilde{T}(\varepsilon,\delta)}\dots\sum_{t_{S}/6=0}^{\tilde{T}(\varepsilon,\delta)}\mathbb{P}(\{T_{i}=t_{i}\}_{1\leq i\leq S/6})\mathbb{P}(\{\Lambda_{i}^{c}\}_{1\leq i\leq S/6}\cap\{T_{i}=t_{i}\}_{1\leq i\leq S/6})$
$=\sum_{t_{1}=0}^{\tilde{T}(\varepsilon,\delta)}\dots\sum_{t_{S}/6=0}^{\tilde{T}(\varepsilon,\delta)}\mathbb{P}(\{T_{i}=t_{i}\}_{1\leq i\leq S/6})\prod_{1\leq i\leq S/6}\mathbb{P}(\Lambda_{i}^{c}\cap\{T_{i}=t_{i}\})$
$=\sum_{t_{1}=0}^{\tilde{T}(\varepsilon,\delta)}\dots\sum_{t_{S}/6=0}^{\tilde{T}(\varepsilon,\delta)}\mathbb{P}(\{T_{i}=t_{i}\}_{1\leq i\leq S/6})(1-\delta)^{S/6}\,,$
@@
($\mathbb{P}\big{(}\{\Lambda_{i}^{c}\}_{1\leq i\leq S/6}|\{T_{i}\leq\tilde{T}(\varepsilon,\delta)\}_{1\leq i\leq S/6}\big{)}\leq(1-\delta)^{\frac{n}{2}}\,.$
Thus, if the total number of transitions $T$ is less than $\frac{S}{6}\tilde{T}(\varepsilon,\delta)$ on the MDP $M_{0}$ corresponding to the hypothesis $H_{0}:\{H_{0}^{i}|1\leq i\leq n\}$, then on $M_{1}$ it holds that
$\mathbb{P}(\|V^{*}-V^{\pi^{\mathcal{U}}_{T}}\|>\varepsilon)$ $\geq\mathbb{P}\bigg{(}\bigcup_{1\leq i\leq S/6}\Lambda(z)\bigg{)}$
$=1-\mathbb{P}\bigg{(}\bigcap_{1\leq i\leq S/6}\Lambda_{i}^{c}\bigg{)}$
$\geq 1-\mathbb{P}(\{\Lambda_{i}^{c}\}_{1\leq i\leq S/6}|\{T_{z_{i}}\leq\tilde{T}(\varepsilon,\delta)\}_{1\leq i\leq S/6})$
$\geq 1-(1-\delta)^{S/6}$
$\geq\frac{\delta S}{12},$
when $\frac{\delta S}{6}\leq 1$ by Lemma 7. By setting $\delta^{\prime}=\delta\frac{S}{12}$ we obtain the result. This shows that if the number of samples is smaller than
$T=\frac{SA}{4800}\log(\frac{S}{96\delta})\frac{\gamma^{2}}{\varepsilon^{2}}\cdot\frac{e^{|\beta|\frac{1}{1-\gamma}}-3}{|\beta|^{2}}$

### v2

@@ H.1 Lower Bound for Q-value Learning
when $\delta\frac{n}{2}\leq 1$ by Lemma 6. By setting $\delta^{\prime}=\delta\frac{n}{4}$ and substituting back $S^{\prime}$ we obtain the result. This shows that if the number of samples is smaller than
$T=\frac{(S^{\prime}-2)A}{4096}\frac{\gamma^{2}}{\varepsilon^{2}}\frac{e^{|\beta|\frac{1}{1-\gamma}}-3}{|\beta|^{2}}\log(\frac{(S^{\prime}-2)A}{32\delta})$ (71)
on the MDP corresponding to the hypothesis $H_{0}:\{H_{0}^{i}|1\leq i\leq n\}$ it holds that $\mathbb{P}_{1}(\|Q^{*}_{M_{1}}-Q^{\mathcal{U}}_{T}\|>\varepsilon)>\delta^{\prime}$.
H.2 Lower Bound for Policy Learning
For a lower bound we construct the following class of MDPs with $S^{\prime}:=S+2$ states and $A^{\prime}:=A+1$ actions where the first states are labelled $s_{1},\ldots,s_{S},s^{G},s^{B}$ and the actions are labelled $a_{0},a_{1},...,a_{A}$. The states $s^{G}$ and $s^{B}$ are absorbing under any actions and $R(s^{G},a)=1$ for all $j$ and $R(s^{B},a)=0$ for all $a\in A$. For the states $s\in\{s_{1},...,s_{S}\}$, we have that $R(s,a)=0$ for all $a\in A$.
From the state $s_{i}$ with probabilities that depend on the action taken the agent will then end up in either a good state $s^{G}$ which is absorbing and yields the maximal unit reward under all actions or in the bad state $s^{B}$ which is also absorbing but which yields no reward under any action. The different MDPs thus differ only in their transition probabilities in the choice states $s_{i}$.
Fix an index $1\leq i\leq S$. We then consider the following set of possible parameters called hypotheses $H^{i}_{l},l\in\{0,1,2,\ldots,A\}$ given by
$H^{i}_{0}:$ $q(s_{i},a_{0})=p+\alpha$ $q(s_{i},a)=p\text{ for }a\neq a_{0}$
$H_{l}:$ $q(s_{i},a_{0})=p+\alpha$ $q(s_{i},a)=p\text{ for }a\notin\{a_{0},l\}\quad$ $q(s_{i},a_{l})=p+2\alpha\,$,
@@ H.2 Lower Bound for Policy Learning
$p$ $=\begin{cases}1-e^{-\beta\frac{1}{1-\gamma}}\quad&\beta>0,\\ e^{-|\beta|\frac{1}{1-\gamma}}\quad&\beta<0,\end{cases}$
@@
Consider a fixed hypothesis $H_{l}^{i}$ for some $l\neq 0$ and the sub-MDP that only consists of the states $\{s_{i},s^{G},s^{B}\}$. Here the optimal action is $a^{*}=a_{l}$, the second best action is $a_{0}$ and all other actions are even worse so the value-error over all states in the triplet for any suboptimal choice of actions will be at least as large as $V^{*}(s_{i})-V^{0}(s_{i})$ where $V^{0}$ is the value by choosing $a=0$. We now show that any non-optimal action is $\varepsilon$-bad on $s_{i}$.
@@
$V^{*}(s_{i})-V^{0}(s_{i})$ $=-\frac{\gamma}{\beta}\log\bigg(\frac{(p+2\alpha)e^{-\beta\frac{1}{1-\gamma}}+1-p-2\alpha}{(p+\alpha)e^{-\beta\frac{1}{1-\gamma}}+1-p-\alpha}\bigg)$
@@
$V^{*}(s_{i})-V^{0}(s_{i})$ $=\frac{\gamma}{|\beta|}\log\bigg(\frac{(p+2\alpha)e^{|\beta|\frac{1}{1-\gamma}}+1-p-2\alpha}{(p+\alpha)e^{|\beta|\frac{1}{1-\gamma}}+1-p-\alpha}\bigg)$
@@
Let $\mathbb{P}_{l}$ and $\mathbb{E}_{l}$ denote the probability operator and expectation operator under the hypothesis $H^{i}_{l}$. Let $t:=t^{i}_{l}$ be the number of times the algorithm tries action $l$ in $s_{i}$ under $H_{0}$. Assuming that $\delta\in(0,\frac{1}{4})$ and using that the algorithm is $(\varepsilon,\delta)$-correct we have that $\mathbb{P}_{0}(B)\geq 1-\delta\geq\frac{3}{4}$ where $B=\{\pi^{\mathcal{U}}(s_{i})=a_{0}\}$ is the event that the algorithm outputs the action $a_{0}$.
@@
Finally, we define the event $\mathcal{E}$ as
$\mathcal{E}=\bigg\{\tilde{p}t-\tilde{k}\leq\sqrt{2p(1-p)\log(\frac{8}{2\theta})}\bigg\}\,.$ (72)
Form the Chernoff-Hoeffding bound and as shown in [23] we have that $\mathbb{P}_{0}(\mathcal{E)}>\frac{3}{4}$ and so $\mathbb{P}_{0}(B\cap\mathcal{E)}>\frac{1}{2}$. From Theorem 9, we get that
$\mathbb{P}_{1}(B)\geq\mathbb{P}_{1}(B\cap\mathcal{E)}=\mathbb{E}_{1}[1_{B}1_{\mathcal{E}}]\geq\mathbb{E}_{0}\bigg[\frac{L_{1}(W)}{L_{0}(W)}1_{\mathcal{E}}1_{B}\bigg]\geq\mathbb{E}_{0}\bigg[\frac{\theta}{4}1_{\mathcal{E}}1_{B}\bigg]=\frac{\theta}{4}\mathbb{P}_{0}(\mathcal{E}\cap B)\geq\frac{\theta}{8}\,.$ (73)
Now solving for $\frac{\theta}{8}>\delta$, we see that if
@@
Since this holds for all the $A$ hypotheses $H_{l}^{i},l=1,2,...,A$ it follows that the algorithm needs at least $\tilde{T}(\varepsilon,\delta):=A\tilde{t}(\varepsilon,\delta)$ samples to be $(\varepsilon,\delta)$-correct on the state $s_{i}$.
Next we use the fact that the structure of the MDPs is such that information used to determine $\pi^{*}(s_{i})$ carries no information to determine $\pi^{*}(s_{j})$ for $i\neq j$.
[…]
Let $T_{i}$ be the number of times the algorithm has tried sampled any action on $s_{i}$ for $i\leq S/2$ Due to the structure of the MDPs in $\mathbb{M}$ it is sufficient to consider only the algorithms that yields an estimate of $\pi^{\mathcal{U}}_{T_{i}}$ based on samples from $s_{i}$ since any other samples can yield no information on $\pi^{*}(s_{i})$.
Thus, by defining the events $\Lambda_{i}:=\{|V_{M_{1}}^{*}(s_{i})-V^{\pi^{\mathcal{U}}_{T_{i}}}(s_{i})|>\varepsilon\}$ we have that $\Lambda_{i}$ and $\Lambda_{j}$ are conditionally independent given $T_{i}$ and $T_{j}$. We then have that for the MDP $M_{1}\in\mathbb{M}$ (The one corresponding to the hypothesis $H_{1}:=\{H_{1}^{i}|1\leq i\leq n\}$) it holds that
$\mathbb{P}(\{\Lambda_{i}^{c}\}_{1\leq i\leq S/2}\cap\{T_{i}\leq\tilde{T}(\varepsilon,\delta)\}_{1\leq i\leq S/2})$
$=\sum_{t_{1}=0}^{\tilde{T}(\varepsilon,\delta)}\dots\sum_{t_{S/2}=0}^{\tilde{T}(\varepsilon,\delta)}\mathbb{P}(\{T_{i}=t_{i}\}_{1\leq i\leq S/2})\mathbb{P}(\{\Lambda_{i}^{c}\}_{1\leq i\leq S/2}\cap\{T_{i}=t_{i}\}_{1\leq i\leq S/2})$
$=\sum_{t_{1}=0}^{\tilde{T}(\varepsilon,\delta)}\dots\sum_{t_{S/2}=0}^{\tilde{T}(\varepsilon,\delta)}\mathbb{P}(\{T_{i}=t_{i}\}_{1\leq i\leq S/2})\prod_{1\leq i\leq S/2}\mathbb{P}(\Lambda_{i}^{c}\cap\{T_{i}=t_{i}\})$
$=\sum_{t_{1}=0}^{\tilde{T}(\varepsilon,\delta)}\dots\sum_{t_{S/2}=0}^{\tilde{T}(\varepsilon,\delta)}\mathbb{P}(\{T_{i}=t_{i}\}_{1\leq i\leq S/2})(1-\delta)^{S/2}\,$,
@@
$\mathbb{P}\big(\{\Lambda_{i}^{c}\}_{1\leq i\leq S/2}|\{T_{i}\leq\tilde{T}(\varepsilon,\delta)\}_{1\leq i\leq S/2}\big)\leq(1-\delta)^{\frac{S}{2}}\,.$
Thus, if the total number of transitions $T$ is less than $\frac{S}{2}\tilde{T}(\varepsilon,\delta)$ on the MDP $M_{0}$ corresponding to the hypothesis $H_{0}:\{H_{0}^{i}|1\leq i\leq n\}$, then on $M_{1}$ it holds that
$\mathbb{P}(\|V^{*}-V^{\pi^{\mathcal{U}}_{T}}\|>\varepsilon)$ $\geq\mathbb{P}\bigg(\bigcup_{1\leq i\leq S/2}\Lambda(z)\bigg)$
$=1-\mathbb{P}\bigg(\bigcap_{1\leq i\leq S/2}\Lambda_{i}^{c}\bigg)$
$\geq 1-\mathbb{P}(\{\Lambda_{i}^{c}\}_{1\leq i\leq S/2}|\{T_{z_{i}}\leq\tilde{T}(\varepsilon,\delta)\}_{1\leq i\leq S/2})$
$\geq 1-(1-\delta)^{S/2}$
$\geq\frac{\delta S}{4},$
when $\frac{\delta S}{2}\leq 1$ by Lemma 6. By setting $\delta^{\prime}=\delta\frac{S}{4}$ and substituting back $S^{\prime}$ and $A^{\prime}$ we obtain the result. This shows that if the number of samples is smaller than
$T=\frac{(S^{\prime}-2)(A^{\prime}-1)}{1600}\log(\frac{S^{\prime}-2}{32\delta})\frac{\gamma^{2}}{\varepsilon^{2}}\cdot\frac{e^{|\beta|\frac{1}{1-\gamma}}-3}{|\beta|^{2}}$